OpenStudy (anonymous):

How many ways can four runners finish a race? A. 10 B. 24 C. 4 D. 1 is it 10?

OpenStudy (anonymous):

no it is not ten

OpenStudy (anonymous):

4! probably

OpenStudy (anonymous):

4 choices for first place 3 for second 2 for third 1 for last by the counting principle the number of ways this can be done is \[4\times 3\times 2\times 1\]

OpenStudy (pinksapphire):

The answer is B.

OpenStudy (pinksapphire):

Try the counting principle.

OpenStudy (tkhunny):

Running Walking Crawling Rolling Wheel-Barrowing Carried Skipping ... I could go on. I do wish we had better questions in out textbooks!

OpenStudy (pinksapphire):

Is the question asking for what places the runners can finish in?

OpenStudy (tkhunny):

Assuming there are no ties and all four runners finish the race, how many different positions can there be? Please read satellite73's explanation.

OpenStudy (anonymous):

its 4 factorial 4! =24